# Fast, Stable and Efficient Approximation of Multi-parameter Persistence Modules with MMA

### v3

@@ 4.2 MMA algorithm for computing candidate decompositions
    match together bars that correspond to the same underlying summand using the matching function $\sigma$,
@@
Step 1 can be performed using any persistent homology software (such as, e.g., Gudhi, Ripser, Phat, etc), or with Rivet [LW15] when $n=2$. Our code can be found at https://github.com/DavidLapous/multipers, and is based on the vineyard algorithm [CSEM06], which allows us to run Steps 1 and 2 jointly (see Section 5.2).
Input 1: A multi-parameter persistence module ${\mathbb{M}}$ (whose multigraded Betti numbers are included in a hyperrectangle $K$),
@@
Input 3: Matching function $\sigma$ (commuting with $\sigma_{{\mathbb{M}}}$ if ${\mathbb{M}}$ is interval decomposable)
Output: Candidate decomposition $\tilde{\mathbb{M}}_{\delta}$
Compute ${\mathcal{F}}{\mathcal{B}}({\mathbb{M}})_{L}={\mathcal{F}}{\mathcal{B}}(\mathcal{M})_{L}$, i.e., the $L$-fibered barcode of ${\mathbb{M}}$;
$S\leftarrow$ []; # $S$ is the set of interval summands of the output candidate decomposition, intialized as the empty set
@@
       for $[b_{l}^{\mathbb{M}},d_{l}^{\mathbb{M}}]\in{\mathcal{B}}({\left.\kern-1.2pt{\mathbb{M}}\vphantom{\big{|}}\right|_{l}})$ do
@@
             if $\exists B\in S$ and $[b,d]\in B$ s.t. $[b_{l}^{{\mathbb{M}}},d_{l}^{{\mathbb{M}}}]=\sigma([b,d])$ then
                   $B$.append($[b_{l}^{{\mathbb{M}}},d_{l}^{{\mathbb{M}}}]$); # If it is, attach the bar to the corresponding summand
@@
Return $\tilde{\mathbb{M}}_{\delta}:=\bigoplus_{B\in S}$ ApproximateInterval$(B)$;
@@
Input: Set of bars $B=\{[b_{l},d_{l}]\}_{l\in L_{B}}$, where $L_{B}\subseteq L$
Output: Discretely presented interval module $\tilde{I}$
labs $\leftarrow$ LabelEndpoints$(B)$;
@@
Return $\tilde{I}(B):=\mathrm{\mathbf{Int}}\left(\bigcup_{c\in C_{B}^{L}(I)}\bigcup_{c^{\prime}\in C_{D}^{L}(I)}R_{c,c^{\prime}}\right)$;
Algorithm 2 ApproximateInterval
We first describe LabelEndpoints. The core idea of this algorithm, whose pseudo-code is given in Algorithm 3, is, for a given bar in $I$ associated to a line $l\in L$, to look at the corresponding surrounding set $L_{l}$. If there exists a hyperplane $H$ such that all endpoints in this surrounding set belong to $H$, we identify $H$ as a facet, and we label the bar with the codirection of $H$.
Input: Set of bars $B=\{[b_{l},d_{l}]\}_{l\in L_{B}}$, where $L_{B}\subseteq L$
Output: List labs of labels for each endpoint in $B$
labs$(b_{l})\leftarrow[]$ for all $l\in L_{B}$;
@@
Figure 7: Example of birthpoint labelling for an interval module $I$ with rectangle support with three surrounding sets of lines $L_{l_{1}}$, $L_{l_{2}}$, $L_{l_{3}}$ associated to three lines $l_{1},l_{2},l_{3}$. The labels of $l_{1},l_{2},l_{3}$ that are identified correspond to the red, blue and grey colored facets of $I$ respectively.
Finally, we describe ComputeCorners. The core idea of the algorithm, whose pseudo-code is given in Algorithm 4, is to use the labels identified by LabelEndpoints to compute corners, or critical point estimates, in the following way: if all birthpoints (resp. deathpoints) in a surrounding set have at least one associated facet, i.e., have a non-empty list of labels, then a candidate corner can be defined using the minimum (resp. maximum) of all birthpoints (resp. deathpoints) coordinates. We only present the pseudo-code for birthpoints since the code for deathpoints is symmetric and can be obtained by replacing minimum by maximum and $\infty$- ∞ by +$\infty$+ ∞.
@@
             if labs$(b)\neq\varnothing,\forall b\in B_{L_{l}}$ then
                   $\{(j,c_{j})\,:\,j\in\mathcal{J}\}\leftarrow\bigcup_{b\in B_{L_{l}}}$ labs$(b)$; # $\mathcal{J}\subseteq\llbracket 1,n\rrbracket$ is the set of codirections
@@
             Assert labs$(b)\neq\varnothing$ for all $b\in B_{L_{l}}$;
             $\{(j,c_{j})\,:\,j\in\mathcal{J}\}\leftarrow\bigcup_{b\in B_{L_{l}}}\textsc{labs}(b)$; # The cardinality of the set of codirections $\mathcal{J}\subsetneq\llbracket 1,n\rrbracket$ must be strictly less than $n$
             Define $C^{l}\in\mathbb{R}^{n}$ as
@@
Computing the $L$-fibered barcode ${\mathcal{F}}{\mathcal{B}}({\mathbb{M}})_{L}$ on a simplicial complex, as well as assigning the corresponding bars to their associated summands in the decomposition of ${\mathbb{M}}$, can be done with the vineyard algorithm and matching [CSEM06] with complexity $O(N^{3}+|L|\cdot N\cdot T)$, where $N$ is the number of simplices in the simplicial complex, and $T$ is the maximal number of transpositions required to update the single-parameter filtrations corresponding to the consecutive lines in $L$. In the worst case scenario, $T=N^{2}$. Note that $T$ usually decreases to a fixed constant as $|L|$ increases, and that this computation can be easily parallelized in practice.
Now, adding the complexities of Algorithms 3 and 4, the final complexity of Algorithm 1 is
@@
Of importance, the dependence on $n$ is much better than the (exact) decomposition algorithm proposed in [DX22] whose complexity is $O(N^{n(2\omega+1)})$, where $\omega<2.373$ is the matrix multiplication exponent. It is also comparable to Rivet [LW15] (which works only when $n=2$), whose complexity is $O(N^{3}\kappa+(N+{\rm log}\kappa)\kappa^{2})$, where $\kappa=\kappa_{x}\kappa_{y}$ is the product of $x$ and $y$ coordinates used to evaluate the module (note that $\kappa_{x},\kappa_{y}$ are also user-dependent). The elder-rule staircode [CKMW20] works only for point cloud data when $n=2$ and homology dimension 00, but has better complexity $O(m^{2}\log(m))$, where $m$ is the number of points. Finally, note that our complexity can be controlled by the number of lines, which is user-dependent. We illustrate this useful property in Section 6.
Remark 4.2.
@@
For the sake of simplicity and efficiency, the code that we provide at https://github.com/DavidLapous/multipers contains a simpler version of Algorithm 4, that does not compute and use labels, but rather gathers the birthpoints and deathpoints as corners directly. One can easily check that the correctness result (Proposition 4.3 and its proof below) carries over to that simpler algorithm, however the exactness result (Corollary 4.10 below) is only valid for corners computed with Algorithm 4.
Correctness of MMA.
The following proposition ensures that MMA is correct.
Proposition 4.3.
Let ${\mathbb{M}}$ be an interval decomposable multi-parameter persistence module. Let $\tilde{\mathbb{M}}^{\mathtt{MMA}{}}_{\delta}=\mathtt{MMA}{}({\mathbb{M}},L_{\delta}(K^{\delta}),\sigma)$, where $\sigma$ commutes with $\sigma_{\mathbb{M}}$. Then $\tilde{\mathbb{M}}^{\mathtt{MMA}{}}_{\delta}$ is a $\delta$-candidate decomposition of ${\mathbb{M}}$.
@@
Let $I$ be an interval summand of ${\mathbb{M}}$, and $\tilde{I}$ be the associated interval summand of the candidate decomposition $\tilde{\mathbb{M}}^{\mathtt{MMA}{}}_{\delta}$ (which is well-defined as $\sigma$ is assumed to commute with $\sigma_{\mathbb{M}}$). Let $C_{B}^{L}(I)$ and $C_{D}^{L}(I)$ be the birth and death corners computed by Algorithm 4, i.e., one has
$\tilde{I}=\mathrm{\mathbf{Int}}\left(\bigcup_{c\in C_{B}^{L}(I)}\bigcup_{c^{\prime}\in C_{D}^{L}(I)}R_{c,c^{\prime}}\right),$ (9)
In order to show that $\tilde{\mathbb{M}}^{\mathtt{MMA}{}}_{\delta}$ is a $\delta$-candidate decomposition of ${\mathbb{M}}$, it suffices to show that the $L$-fibered barcodes of $\tilde{I}$ and $I$ are the same, i.e., ${\mathcal{F}}{\mathcal{B}}(I)_{L}={\mathcal{F}}{\mathcal{B}}(\tilde{I})_{L}$. Equivalently, we need to show that ${\rm supp}(I)\cap l={\rm supp}(\tilde{I})\cap l$ for any line $l\in L$. We first show that they share the same birthpoints, i.e., that $L[I]\cap L=L[\tilde{I}]\cap L$. Let $l\in L$. Note that $b^{I}_{l}$ and $b^{\tilde{I}}_{l}$ are comparable since they belong to the same diagonal line $l$.
Strategy. In order to show $b^{\tilde{I}}_{l}=b^{I}_{l}$, we are going to show that 1. $b^{I}_{l}\leq_{n}b^{\tilde{I}}_{l}$ and 2. $b^{\tilde{I}}_{l}\leq_{n}b^{I}_{l}$.
@@
    In order to show $b^{I}_{l}\leq_{n}b^{\tilde{I}}_{l}$, we are going to show that $c\not<_{n}b^{I}_{l}$ for any corner $c\in C_{B}^{L}(I)$. Indeed, if one assumes $b^{I}_{l}>_{n}b^{\tilde{I}}_{l}$, and since there always exists a birth corner $c\in C_{B}^{L}(I)$ such that $c\leq_{n}b^{\tilde{I}}_{l}$ by construction of $\tilde{I}$, one has $c\leq_{n}b^{\tilde{I}}_{l}<_{n}b^{I}_{l}$.
@@
    In order to show $b^{\tilde{I}}_{l}\leq_{n}b^{I}_{l}$, we are going to show that there exists a corner $c\in C_{B}^{L}(I)$ such that $c\leq_{n}b^{I}_{l}$. Indeed, if there is such a birth corner, and if $b^{\tilde{I}}_{l}>_{n}b^{I}_{l}$ by contradiction, then $c\leq_{n}b^{I}_{l}<_{n}b^{\tilde{I}}_{l}$, and $R_{c,b^{\tilde{I}}_{l}}$ is not flat, contradicting Lemma 3.9.
Proof of (2). By construction of $\tilde{I}$ with Algorithm 4, if $b^{I}_{l}$ is labelled, then there exists a line $l^{\prime}$ and a corner $c^{l^{\prime}}\in C_{B}^{L}(I)$ that is smaller than $b^{I}_{l}$ so we can take $c:=c^{l^{\prime}}$. If $b^{I}_{l}$ is not labelled, it belongs itself to $C_{B}^{L}(I)$, and we can take $c:=b^{I}_{l}$.
Proof of (1). Let $c\in C_{B}^{L}(I)$ be a birth corner, and let $L_{l_{0}}$ be the associated surrounding set of lines for some $l_{0}\in L$. Let $[c]_{l}:=\min\left[\left(c+\left(\mathbb{R}_{+}\right)^{n}\right)\cap l\right]$ be the smallest element in the intersection between the positive cone on $c$ and $l$. Assume $[c]_{l}\geq_{n}b^{I}_{l}$ and $c<_{n}b^{I}_{l}$. Then $R_{c,[c]_{l}}$ is not flat, contradicting the fact that $[c]_{l}$ is the smallest element. Thus, we only have to show $[c]_{l}\geq_{n}b^{I}_{l}$. There are two cases.
  • Either some birthpoints of $L_{l_{0}}$ are not labelled by Algorithm 3, and $c$ is equal to $b^{I}_{l^{\prime}}$ for some $l^{\prime}\in L_{l_{0}}$. Now, assume $[c]_{l}<_{n}b^{I}_{l}$ by contradiction. Then $b^{I}_{l^{\prime}}=c\leq_{n}[c]_{l}<_{n}b^{I}_{l}$. Thus $b^{I}_{l^{\prime}}<_{n}b^{I}_{l}$ and $R_{b^{I}_{l^{\prime}},b^{I}_{l}}$ is not flat, contradicting Lemma 3.9. Hence $[c]_{l}\geq_{n}b^{I}_{l}$.
@@
    • Either $l\in L_{l_{0}}$. Then, $\exists i\in\llbracket 1,n\rrbracket$ such that $(b^{I}_{l})_{i}=c_{i}$. This yields $(b^{I}_{l})_{i}=c_{i}\leq([c]_{l})_{i}$, and thus $[c]_{l}\geq_{n}b^{I}_{l}$ since they both belong to the same diagonal line $l$.
    • Or the line $l$ does not belong to $L_{l_{0}}$. Since $[c]_{l}$ is on the boundary of the positive cone based on $c$, there exists $i\in\llbracket 1,n\rrbracket$ such that $([c]_{l})_{i}=c_{i}$. Assume again by contradiction that $b^{I}_{l}>_{n}[c]_{l}$, and write
      $[c]_{l}=c+\sum_{j\neq i}(\delta\alpha_{j})e_{j}=:c+\overrightarrow{v}<_{n}b^{I}_{l}$
      with $\alpha_{j}\geq 0$ for $j\in\llbracket 1,n\rrbracket\backslash\left\{i\right\}$. Since $l\notin L_{l_{0}}$, there exists some $j_{0}$ such that $\alpha_{j_{0}}>1$. Let $\overrightarrow{u}:=((\overrightarrow{v}_{j}\,\mathrm{mod}\,\delta)_{j\in\llbracket 1,n\rrbracket})=(([c]_{l}-c)_{j}\,\mathrm{mod}\,\delta)_{j\in\llbracket 1,n\rrbracket}\in[0,\delta)^{n}\leq_{n}\overrightarrow{v}$. Let $l^{\prime}:=l_{c+\overrightarrow{u}}$ be the diagonal line passing through $c+\overrightarrow{u}$. Now, recall that the lines of $L$ are drawn on a grid, so $l^{\prime}\in L$ since $l^{\prime}=l+\overrightarrow{u}-\overrightarrow{v}$. Moreover, one has by definition, $c\in{\rm conv}(L_{l_{0}})$. Since the lines of $L$ are on a grid, one has
      $\forall l_{1},l_{2}\in L,\|l_{1}\cap H_{n},{\rm conv}(L_{l_{2}})\cap H_{n})\|_{\infty}<\delta\Longrightarrow l_{1}\in L_{l_{2}},\quad$
      where $H_{n}=\left\{x\in\mathbb{R}^{n}:x_{n}=c_{n}\right\}$. Now, note that $c+\overrightarrow{u}$ and $c+\overrightarrow{u}-\overrightarrow{u}_{n}\cdot\textbf{1}$ both belong to $l^{\prime}$, and that $c+\overrightarrow{u}-\overrightarrow{u}_{n}\cdot\textbf{1}\in H_{n}$. Moreover, since

### v4

@@ 3.3 The MMA algorithm for computing candidate decompositions
    match together bars using the matching function $\sigma$,
@@
Step 1 can be performed using any persistent homology software (such as, e.g., Gudhi, Ripser, Phat, etc), or with Rivet (47) when $n=2$. Our code is part of the multipers library (50), and can be found at https://github.com/DavidLapous/multipers. Moreover, it uses the vineyard algorithm (28), which allows us to run Steps 1 and 2 jointly (see Section 6.2).
Input 1: A f.p. $n$-parameter persistence module $\mathbb{M}$} and a compact $K\subset\mathbb{R}^{n}$ that compactly characterizes $\mathbb{M}$,
@@
Input 3: Matching function $\sigma$
Output: Candidate decomposition $\tilde{\mathbb{M}}^{\texttt{MMA}{}}_{\delta}$
Compute ${\mathcal{F}}{\mathcal{B}}(\mathbb{M})_{L}$, i.e., the $L$-fibered barcode of $\mathbb{M}$;
$\leftarrow []; # SS$ is the set of interval summands of the output candidate decomposition, intialized as the empty set
@@
 for $b_{l}^{\mathbb{M}},d_{l}^{\mathbb{M}}]\in{\mathcal{B}}\left({\left.\kern-1.2pt\mathbb{M}\vphantom{\big|}\right|_{l}}\right)$ do
@@
    if $\exists B\in S$ and $b,d]\in B$ s.t. $b_{l}^{\mathbb{M}},d_{l}^{\mathbb{M}}]=\sigma([b,d])$ then
       $B$.append($b_{l}^{\mathbb{M}},d_{l}^{\mathbb{M}}]$); # If it is, attach the bar to the corresponding summand
@@
Return $\tilde{\mathbb{M}}^{\texttt{MMA}{}}_{\delta}:=\bigoplus_{B\in S}$ ApproximateInterval(B)(B);
@@
Input: Set of bars $B=\{[b_{l},d_{l})\}_{l\in L_{B}}$, where $L_{B}\subseteq L$
Output: Discretely presented interval module $k^{\tilde{I}(B)}$}
labs $\leftarrow$ LabelEndpoints(B)(B);
$C_{B}^{L}(B),C_{D}^{L}(B)\leftarrow$ ComputeCorners(B,(B, labs));
$\tilde{I}(B)$ $\leftarrow$ $\bigcup_{c\in C_{B}^{L}(B)}\bigcup_{c^{\prime}\in C_{D}^{L}(B)}R_{c,c^{\prime}}$;
Return $k^{\tilde{I}(B)}$;
Algorithm 2 ApproximateInterval
We first describe LabelEndpoints. The core idea of this algorithm, whose pseudo-code is given in Algorithm 3, is, for a given bar in $I$ associated to a line $l\in L$, to look at the corresponding surrounding set $L_{l}$ (see item (4) in Definition 17). If there exists a hyperplane $H$ such that all endpoints in this surrounding set belong to $H$, we identify $H$ as a facet, and we label the bar with the codirection of $H$.
Input: Set of bars $B=\{[b_{l},d_{l})\}_{l\in L_{B}}$, where $L_{B}\subseteq L$
Output: List labs of labels for each endpoint in $B$
@@
Figure 6: Example of birthpoint labelling for an interval module $I$ with rectangle support with three surrounding sets of lines $L_{l_{1}}$, $L_{l_{2}}$, $L_{l_{3}}$ associated to three lines $l_{1},l_{2},l_{3}$. The labels of $l_{1},l_{2},l_{3}$ that are identified correspond to the red, blue and grey colored facets of II respectively.
Finally, we describe ComputeCorners. The core idea of the algorithm, whose pseudo-code is given in Algorithm 4, is to use the labels identified by LabelEndpoints to compute corners, or critical point estimates, in the following way: if all birthpoints (resp. deathpoints) in a surrounding set have at least one associated facet, i.e., have a non-empty list of labels, then a candidate corner can be defined using the minimum (resp. maximum) of all birthpoints (resp. deathpoints) coordinates. We only present the pseudo-code for birthpoints since the code for deathpoints is symmetric and can be obtained by replacing minimum by maximum and -$\infty$ by $\infty$. Note that the correctness of MMA follows directly from how these corners are computed: it is clear from Algorithm 4 that, for any bar in the fibered barcode, the algorithm will produce a corner that is lower (resp. larger) w.r.t. the partial order $\leq_{n}$ than the birthpoint (resp. deathpoint) of the bar (excluding the trivial case of corners with infinite coordinates).
@@
[…]
       $\{(j,c_{j})\,:\,j\in\mathcal{J}\}\leftarrow\bigcup_{b\in B_{L_{l}}}$ labs(b)(b); # $\mathcal{J}\subseteq\llbracket 1,n\rrbracket$ is the set of codirections
@@
[…]
    $\{(j,c_{j})\,:\,j\in\mathcal{J}\}\leftarrow\bigcup_{b\in B_{L_{l}}}\textsc{labs}(b)$; # The cardinality of $\mathcal{J}$ must be strictly less than $n$
      Define $C^{l}\in\mathbb{R}^{n}$ as:
@@
Computing the $L$-fibered barcode ${\mathcal{F}}{\mathcal{B}}(\mathbb{M})_{L}$ on a simplicial complex, as well as assigning the corresponding bars to their associated summands in the decomposition of $\mathbb{M}$, can be done with the vineyard algorithm (28) as matching function with complexity $O(N^{3}+|L|\cdot N\cdot T)$, where $N$ is the number of simplices in the simplicial complex, and $T$ is the maximal number of transpositions required to update the single-parameter filtrations corresponding to the consecutive lines in $L$. In the worst-case scenario, one has $T=N^{2}$. Note that $T$ usually decreases to a fixed constant as $|L|$ increases, and that this computation can be easily parallelized in practice.
Now, adding the complexities of Algorithms 3 and 4, the final complexity of Algorithm 1 is:
@@
Of importance, the dependence on $n$ is much better than the (exact) decomposition algorithm proposed in (34, 32) whose complexity is $O(N^{n(2\omega+1)})$, where $\omega<2.373$ is the matrix multiplication exponent. It is also comparable to Rivet (47) (although Rivet only works when $n=2$), whose complexity is $O(N^{3}\kappa+(N+{\rm log}\kappa)\kappa^{2})$, where $\kappa=\kappa_{x}\kappa_{y}$ is the product of $x$ and $y$ coordinates used to evaluate the module (note that $\kappa_{x},\kappa_{y}$ are also user-dependent). The elder-rule staircode (17) works only for point cloud data when $n=2$ and homology dimension 0, but has better complexity $O(m^{2}\log(m))$, where $m$ is the number of points. Finally, note that our complexity can be controlled by the number of lines, which is user-dependent. We illustrate this useful property in Section 7.
Remark 8.
@@
For the sake of simplicity and efficiency, the code that we provide at https://github.com/DavidLapous/multipers contains a simpler version of Algorithm 4, that does not compute and use labels, but rather gathers the birthpoints and deathpoints as corners directly. One can easily check that our approximation guarantees (Proposition 3 and Proposition 5) carry over to that simpler algorithm, however the exactness result (Proposition 12) is only valid for corners computed with Algorithm 4.
4 Theoretical robustness of MMA
In this section, our goal is to prove our first important result, Proposition 3, which states that, if the matching function is compatible, then the candidate decompositions computed by MMA are also approximate decompositions: they preserve the diagonal barcodes (up to $2\delta$) associated to all diagonal lines. We provide a proof in Section 4.1. We also discuss the stability of MMA w.r.t. $d_{\rm I}$} in Section 4.2.
4.1 Approximation guarantee of MMA
We first introduce so-called compatible matching functions, which are key elements for proving the approximation property satisfied by our MMA algorithm.
Definition 22 (Compatible matching function).
Let $\mathbb{M}$ be a f.p. $n$-parameter persistence module, and let $l_{1},l_{2}\subseteq\mathbb{R}^{n}$ be two diagonal lines that are at distance $\delta$ from each other. Assume ${\rm supp}(\mathbb{M})\cap l_{1}$ and ${\rm supp}(\mathbb{M})\cap l_{2}$ are not empty, and let $b_{1},d_{1})$ and $b_{2},d_{2})$ be bars in ${\mathcal{B}}\left({\left.\kern-1.2pt\mathbb{M}\vphantom{\big|}\right|_{l_{1}}}\right)$ and ${\mathcal{B}}\left({\left.\kern-1.2pt\mathbb{M}\vphantom{\big|}\right|_{l_{2}}}\right)$, characterized by their endpoints. These bars are compatible if the rectangles $R_{b_{1},b_{2}}$, $R_{b_{2},b_{1}}$, $R_{d_{1},d_{2}}$ and $R_{d_{2},d_{1}}$ are flat or empty. Equivalently, two bars are compatible if their birthpoints (resp. deathpoints) are not strictly comparable, i.e., $b_{1}\not<_{n}b_{2}$, and $b_{1}\not>_{n}b_{2}$ (resp. $d_{1}\not<_{n}d_{2}$, and $d_{1}\not>_{n}d_{2}$). Moreover, we say that $b^{\mathbb{M}}_{l_{1}},d^{\mathbb{M}}_{l_{1}})$ is compatible with the empty set in $l_{2}$ if $\left\lVert b^{\mathbb{M}}_{l_{1}}-d^{\mathbb{M}}_{l_{1}}\right\rVert_{\infty}\leq 2\delta$.
Remark 9.
Induced matching functions (see Definition 21) are compatible, as per Lemma 1.
Proposition 3 (Approximation result).
Let $\mathbb{M}$ be a f.p. $n$-parameter persistence module and $\delta>0$. Let $K$ be a rectangle in $\mathbb{R}^{n}$ that compactly characterizes $\mathbb{M}$, and $L:=L_{\delta}(K^{\delta})$ be the $\delta$-grid of lines of the offset $K^{\delta}$. Finally, let $\tilde{\mathbb{M}}^{\mathtt{MMA}{}}_{\delta}:=\mathtt{MMA}{}({\mathbb{M}},L,\sigma)$}, where $\sigma$ is a compatible matching function. Then $\tilde{\mathbb{M}}^{\mathtt{MMA}{}}_{\delta}$ is a $\delta$-approximate decomposition of $\mathbb{M}$. More precisely, given some diagonal line $l$, one has:
  1. (i{i})
    ${d_{\rm B}}\left({\mathcal{B}}\left({\left.\kern-1.2pt\mathbb{M}\vphantom{\big|}\right|_{l}}\right),{\mathcal{B}}\left({\left.\kern-1.2pt\tilde{\mathbb{M}}^{\emph{{MMA}{}}}_{\delta}\vphantom{\big|}\right|_{l}}\right)\right)=0$ if $l\in L$, and
  2. (ii{ii})
    ${d_{\rm B}}\left({\mathcal{B}}\left({\left.\kern-1.2pt\mathbb{M}\vphantom{\big|}\right|_{l}}\right),{\mathcal{B}}\left({\left.\kern-1.2pt\tilde{\mathbb{M}}^{\emph{{MMA}{}}}_{\delta}\vphantom{\big|}\right|_{l}}\right)\right)\leq 2\delta$ otherwise.
In order for Proposition 3 to apply, one needs to find a compatible matching function $\sigma$. We discuss how to design such matching functions for interval decomposable modules in Section 6.1 and for general $2$-parameter modules in Section 6.2. We also hypothesize that compatible matching functions for general $n$-parameter persistence modules could be constructed using representative cycles in a similar way than the construction of the graphcode (39, 40), a conjecture that we leave for future work.
@@ 4.1 Approximation guarantee of MMA
We first prove (i{i}). Let $l\in L$, and let $b_{l}$ be the birthpoint of a bar $b$ in ${\mathcal{B}}\left({\left.\kern-1.2pt\mathbb{M}\vphantom{\big|}\right|_{l}}\right)$. Let $B$ be the set of bars containing $b$ computed with Algorithm 1, let $C_{B}^{L}(B)$ and $C_{D}^{L}(B)$ be the birth and death corners computed with Algorithm 4, and let $\tilde{I}$ be the interval computed with Algorithm 2, i.e., one has:
$\tilde{I}=\bigcup_{c\in C_{B}^{L}(B)}\bigcup_{c^{\prime}\in C_{D}^{L}(B)}R_{c,c^{\prime}},$ (10)
In order to show (i{i}), we first need to show that $b_{l}=b^{\tilde{I}}_{l}$ (and then the proof for deathpoints will follow by symmetry), where $b^{\tilde{I}}_{l}$ is defined as per Definition 15. Note that $b_{l}$ and $b^{\tilde{I}}_{l}$ are comparable since they belong to the same diagonal line $l$.
Strategy. In order to show $b_{l}=b^{\tilde{I}}_{l}$, we are going to show that 1. $b_{l}\leq_{n}b^{\tilde{I}}_{l}$ and 2. $b^{\tilde{I}}_{l}\leq_{n}b_{l}$.
@@
    In order to show $b_{l}\leq_{n}b^{\tilde{I}}_{l}$, we are going to show that $c\not<_{n}b_{l}$ for any corner $c\in C_{B}^{L}(B)$. Indeed, if one assumes $b_{l}>_{n}b^{\tilde{I}}_{l}$ by contradiction, and since there always exists a birth corner $c\in C_{B}^{L}(B)$ such that $c\leq_{n}b^{\tilde{I}}_{l}$ by construction of $\tilde{I}$, one has $c\leq_{n}b^{\tilde{I}}_{l}<_{n}b_{l}$.
@@
    In order to show $b^{\tilde{I}}_{l}\leq_{n}b_{l}$, we are going to show that there exists a corner $c\in C_{B}^{L}(B)$ such that $c\leq_{n}b_{l}$. Indeed, if there is such a birth corner, and if $b^{\tilde{I}}_{l}>_{n}b_{l}$ by contradiction, then $c\leq_{n}b_{l}<_{n}b^{\tilde{I}}_{l}$, and $R_{c,b^{\tilde{I}}_{l}}$ is not flat, contradicting Lemma 1.
Proof of (2). By construction of $\tilde{I}$ with Algorithm 4, if $b_{l}$ is labelled, then there exists a line $l^{\prime}$ and a corner $c^{l^{\prime}}\in C_{B}^{L}(B)$ that is smaller than $b_{l}$ so we can take $c:=c^{l^{\prime}}$. If $b_{l}$ is not labelled, it belongs itself to $C_{B}^{L}(B)$, and we can take $c:=b_{l}$.
Proof of (1). Let $c\in C_{B}^{L}(B)$ be a birth corner, and let $L_{l_{0}}$ be the associated surrounding set of lines for some $l_{0}\in L$. Let $c]_{l}:=\min\left[\left(c+\left(\mathbb{R}_{+}\right)^{n}\right)\cap l\right]$ be the smallest element in the intersection between the positive cone on $c$ and $l$. Assume $c]_{l}\geq_{n}b_{l}$ and $c<_{n}b_{l}$. Then $R_{c,[c]_{l}}$ is not flat, contradicting the fact that $c]_{l}$ is the smallest element. Thus, we only have to show $c]_{l}\geq_{n}b_{l}$. There are two cases.
  1. 1.
    Either some birthpoints of $L_{l_{0}}$ are not labelled by Algorithm 3, and $c$ is equal to the birthpoint $b_{l^{\prime}}$ of another bar in ${\mathcal{B}}\left({\left.\kern-1.2pt\mathbb{M}\vphantom{\big|}\right|_{l^{\prime}}}\right){\cap B}$ for some $l^{\prime}\in L_{l_{0}}$. Now, assume $c]_{l}<_{n}b_{l}$ by contradiction. Then $b_{l^{\prime}}=c\leq_{n}[c]_{l}<_{n}b_{l}$. Thus $b_{l^{\prime}}<_{n}b_{l}$ and $R_{b_{l^{\prime}},b_{l}}$ is not flat, contradicting the fact that $\sigma$ is compatible. Hence $c]_{l}\geq_{n}b_{l}$.
  2. 2.
@@
    1. (a)
      Either $l\in L_{l_{0}}$. Then, $\exists i\in\llbracket 1,n\rrbracket$ such that $(b_{l})_{i}=c_{i}$. This yields $(b_{l})_{i}=c_{i}\leq([c]_{l})_{i}$, and thus $c]_{l}\geq_{n}b_{l}$ since they both belong to the same diagonal line $l$.
    2. (b)
      Or the line $l$ does not belong to $L_{l_{0}}$. Since $c]_{l}$ is on the boundary of the positive cone based on $c$, there exists $i\in\llbracket 1,n\rrbracket$ such that $([c]_{l})_{i}=c_{i}$. Assume again by contradiction that $b_{l}>_{n}[c]_{l}$, and write:
      $c]_{l}=c+\sum_{j\neq i}(\delta\alpha_{j})e_{j}=:c+\overrightarrow{v}<_{n}b_{l},$
      with $\alpha_{j}\geq 0$ for $j\in\llbracket 1,n\rrbracket\backslash\left\{i\right\}$. Since $l\notin L_{l_{0}}$, there exists some $j_{0}$ such that $\alpha_{j_{0}}>1$. Let $\overrightarrow{u}:=((\overrightarrow{v}_{j}\,\mathrm{mod}\,\delta)_{j\in\llbracket 1,n\rrbracket})=(([c]_{l}-c)_{j}\,\mathrm{mod}\,\delta)_{j\in\llbracket 1,n\rrbracket}\in[0,\delta)^{n}\leq_{n}\overrightarrow{v}$. Let $l^{\prime}:=l_{c+\overrightarrow{u}}$ be the diagonal line passing through $c+\overrightarrow{u}$. Now, recall that the lines of $L$ are drawn on a grid, so $l^{\prime}\in L$ since $l^{\prime}=l+\overrightarrow{u}-\overrightarrow{v}$. Moreover, one has: by definition, $c\in{\rm conv}(L_{l_{0}})$. Since the lines of $L$ are on a grid, one has:
      $\forall l_{1},l_{2}\in L,\quad\|l_{1}\cap H_{n},{\rm conv}(L_{l_{2}})\cap H_{n})\|_{\infty}<\delta\Longrightarrow l_{1}\in L_{l_{2}},$
      where $H_{n}=\left\{x\in\mathbb{R}^{n}:x_{n}=c_{n}\right\}$. Now, note that $c+\overrightarrow{u}$ and $c+\overrightarrow{u}-\overrightarrow{u}_{n}\cdot\textbf{1}$ both belong to $l^{\prime}$, and that $c+\overrightarrow{u}-\overrightarrow{u}_{n}\cdot\textbf{1}\in H_{n}$. Moreover, since: